# On multi-graded Proj schemes

### v4

@@ 1.1 Proj schemes
As part of his refoundation of Algebraic Geometry during the second part of the twentieth century, A. Grothendieck introduced the Proj scheme of an ${\mathbb{N}}$-graded ring. This generalization of the notion of projective space is natural in the framework of his theory of schemes. Grothendieck’s Proj construction has since been studied, taught, treated and used many thousands of times. During the first years of the third millennium, O. Brenner and S. Schröer generalized Grothendieck’s Proj construction to rings graded by arbitrary finitely generated abelian groups. Surprisingly, this very nice generalization still seems to be relatively confidential, although it has been recently used and studied in some references, including [Za19, KSU21, MRo24]. In this paper, we provide a slightly different perspective on the Brenner–Schröer Proj construction, and prove a couple of new results on this construction. In particular, we explain a generalization in this setting of Serre’s classical description of coherent sheaves on a projective space (or, more generally, a Proj scheme) as a Serre quotient of a category of graded modules over a graded ring, see e.g. [StP, Tag 01YR].
To illustrate this theory, we also explain how to describe flag varieties of reductive algebraic groups over algebraically closed fields, and vector bundles on such flag varieties associated with some representations of a Borel subgroup (including the Springer resolution) as the Proj scheme of a ring graded by characters of a maximal torus. The description of flag varieties as Proj schemes of $\mathbb{N}$-graded rings is classical, see e.g. [Wa], but it requires a choice of a strictly dominant weight. In the multi-graded setting one does not make any such choice, hence gets a more canonical construction, which is sometimes useful. In fact, versions of this construction appear implicitly or explicitly (but without proper references) in various constructions in Geometric Representation Theory, including [AB09, ABG04, ARd16], and it was one of our motivations for this work to make this construction completely explicit and rigorous. Note that [MS05, Exercise 14.16] indicates a special case of this construction.
@@ 1.2 The Brenner–Schröer construction
Let us now introduce the ideas of our work more precisely. Recall that given a commutative ring $A$, its (affine) spectrum ${\operatorname{Spec}}(A)$ is defined as the set of prime ideals, endowed with a certain topology and a sheaf of rings. Most treatments of Grothendieck’s Proj construction proceed by analogy with this construction, namely by associating to a graded ring the set of its graded prime ideals, and endowing it with a topology and a sheaf of rings after (somewhat tedious) identifications of various sets. However, Brenner–Schröer proceed differently in their generalization in [BS03, §2]: they start from the fact that, if $M$ is an abelian group, the datum of a grading on a commutative ring $A$ is equivalent to the datum of an action of the associated diagonalizable group scheme $\mathrm{D}_{{\operatorname{Spec}}(\mathbb{Z})}(M)$ on ${\operatorname{Spec}}(A)$ (see e.g. [SGA3, Exp. I, §4.7.3]), and then consider a certain open subset in a quotient ringed space constructed in terms of this action. Brenner–Schröer’s Proj scheme is covered by some special affine open subschemes, which generalize similar special affine subschemes appearing in Grothendieck’s Proj construction. In case $M=\mathbb{Z}$ and the negative-degree components in $A$ are 00, this recovers Grothendieck’s construction. (We will refer to this case as the “$\mathbb{N}$-graded case.”)
In the present treatment of the Proj construction, we put these “special affine subschemes” at the heart of the Proj construction. In fact, we give them a name: “potions.” Potions glue together to give birth to the Proj scheme thanks to a purely algebraic statement we call “the magic of potions.” To use this statement and glue potion schemes, we have to assume that $M$ is finitely generated. (In the $\mathbb{N}$-graded case, this result is part, as an auxiliary lemma, of the justification of the tedious identifications mentioned before in most treatments of the Proj construction.) Note that, in general, the underlying set of the Proj scheme is not the set of graded prime ideals of the given graded ring; it can be described as a set of graded ideals (see [BS03, Remark 2.3]), but we believe that this point of view is not really useful, and in any case it does not play any role in the present paper. Note that [MS05], already mentioned, provides some specific multi-graded Proj schemes (under the name spector) by glueing, cf. [MS05, Definition 10.25]. It seems that [BS03] and [MS05] are independent (each does not cite the other). Some other works related to multi-graded Proj schemes are cited in [BS03, Introduction].
@@ 1.3 Contents
We now present the results contained in this document. Let as above $M$ be an abelian group and $A$ be a commutative ring endowed with a grading by $M$. In Sections 2 and 3 we define the notions of $M$-relevant families of $A$ (following Brenner–Schröer, see Definition 2.3) and of potions associated with these families, and explain the definition of the Proj scheme ${\operatorname{Proj}}^{M}(A)$ of $A$ (in case $M$ is finitely generated). The potion associated with a relevant family $(f_{i}:i\in I)$ is the degree-zero part in the localization of $A$ with respect to the multiplicative subset generated by $(f_{i}:i\in I)$; see Definition 2.9. The spectra ${\operatorname{Spec}}(A_{(S)})$ are called potion schemes and are denoted $D_{\dagger}(S)$ (cf. Construction 3.1), and the scheme ${\operatorname{Proj}}^{M}(A)$ is defined by glueing the potions schemes $D_{\dagger}(S)$ over all relevant families (cf. Construction 3.1). This glueing is possible by the magic of potions (cf. Proposition 2.10). In other words, Section 2 focuses on commutative algebra around potions while Section 3 applies this material to explain the construction of Proj schemes.
In Sections 2 and 3, we also prove some basic results that may be known to experts. In particular, using the concept of quasi-relevant element (Definition 2.11), we prove that the Proj scheme of a tensor product of graded rings (with the product grading) is the fiber product of the Proj schemes of the rings (Proposition 3.15). We prove some compatibility results regarding the radical of the ideal generated by relevant elements (Propositions 2.5 and 2.7). In Example 2.13, we explain the relation between potions and dilatations of rings (as studied in [MRR23, Ma23b]). We also show that the Proj construction is functorial (Proposition 3.12). In §3.6 we discuss multi-centered blowups and its relations with dilatations.
@@ 1.4 Relations between various constructions for graded rings
Recently, several other basic constructions for ${\mathbb{N}}$-graded rings were extended to more general gradings. In particular, in [Ma23a] the theory of ${\mathbb{G}}_{\mathrm{m}}$-attractors was extended in the setting of actions of diagonalizable group schemes, and in [Ma23b] mono-centered dilatations (involving ${\mathbb{N}}$-gradings) were extended to multi-centered dilatations. The algebraic point of view adopted in the present paper was partly inspired by these parallel generalizations. The diagram in Figure 1 summarizes some relations between these graded constructions. Solid lines materialize immediate connections. If two lines are parallel (not necessarily connected), then the associated connections are also parallel in some appropriate sense.
@@
It is possible that other ${\mathbb{N}}$-graded constructions in schemes theory can be extended significantly to more general gradings.
@@ 1.5 Acknowledgements
We thank R. Bezrukavnikov, S. Brochard, J. P. dos Santos, A. Ducros, and T. Wedhorn for useful discussions and/or comments related to the subject of this paper.
@@ 1.6 Some notations and conventions
1.6.1 Rings
All the rings we will consider in the paper will be tacitly assumed to be associative and unital. (In practice, all the rings we want to consider will be commutative, but we will mention this assumption when it is necessary.)
1.6.2 Graded rings
Let us make explicit our conventions on graded rings. For this we fix an arbitrary abelian group $M$. An $M$-graded ring is a ring $A$ endowed with a direct sum decomposition $A=\bigoplus_{m\in M}A_{m}$ such that $A_{m}\cdot A_{m^{\prime}}\subset A_{m+m^{\prime}}$ for all $m,m^{\prime}\in M$.
Let $A$ be an $M$-graded ring. An element $a\in A$ is called homogeneous if it belongs to $A_{m}$ for some $m\in M$. In this case, if $a$ is nonzero, then the element $m$ is unique; it is denoted $\deg(a)$ and called the degree of $a$.
@@ 1.6.2 Graded rings
The element 00 is homogeneous, but it has no degree. One can easily check that if the unit $1$ is nonzero (i.e. if $A$ is nonzero) then it is homogeneous of degree $0\in M$.
If $R$ is a ring, an $M$-graded $R$-algebra is an $M$-graded ring $A$ endowed with a ring morphism $R\to A_{0}$. (In particular, $A$ is then an $R$-algebra.)
1.6.3 Other conventions
@@ 2.1 Graded rings and localizations
If $A$ is a commutative $M$-graded ring and $S\subset A$ is a homogeneous multiplicative subset, the localization $A_{S}$ of $A$ with respect to $S$ (which is denoted $S^{-1}A$ in many references, including [StP, Tag 00CM]) is canonically $M$-graded, in such a way that $\deg(\frac{a}{s})=\deg(a)-\deg(s)$ for $a\in A$ nonzero homogeneous and $s\in S$ such that $\frac{a}{s}\in A_{S}$ is nonzero. Explicitly, for $m\in M$ we have
$(A_{S})_{m}=\left\{0\right\}\cup\left\{\frac{a}{s}:a\in A\text{ homogeneous nonzero},\,s\in S\text{ such that $\deg(a)-\deg(s)=m$}\right\}.$
Given a graded $A$-module $Q$, one can also consider the associated localization $Q_{S}$ (denoted $S^{-1}Q$ in [StP, Tag 07JZ]), which has a natural structure of graded $A_{S}$-module.
@@ 2.2 Relevant families
Consider an abelian group $M$, and a commutative $M$-graded ring $A$.
@@
Let $S$ be a homogeneous multiplicative subset of $A$. We denote by $\deg(S)$ the subset of $M$ consisting of degrees of nonzero homogeneous elements of $S$.
@@
Remark 2.2.
Note that in general $\deg(S)$ is not a monoid, because a product of nonzero elements of $A$ might be 00.
@@
    A necessary condition for an $M$-relevant homogeneous multiplicative subset to exist is that the degrees of nonzero homogeneous elements of $A$ generate a subgroup $M^{\prime}\subset M$ such that $M/M^{\prime}$ is a torsion abelian group. If this is the case, $0^{\mathbb{N}}$ is an example of such a subset, since $\underline{0^{{\mathbb{N}}}}$ consists of all homogeneous elements of $A$.
@@
Let $M$ and $M^{\prime}$ be two finitely generated abelian groups. Let $R$ be a commutative ring and let $A$ (resp. $A^{\prime}$) be a commutative $M$-graded (resp. $M^{\prime}$-graded) $R$-algebra. Then, considering $A\otimes_{R}A^{\prime}$ with its natural structure of $M\times M^{\prime}$-graded $R$-algebra:
@@
(1) Let $x\in A\otimes_{R}A^{\prime}$ be relevant. Then there exists $k\in\mathbb{Z}_{>0}$ and a factorization $x^{k}=x_{1}\cdots x_{l}$ into homogeneous factors $x_{i}$ of degree $d_{i}\in M\times M^{\prime}$, such that the degrees $d_{i}$ generate a subgroup of $M\times M^{\prime}$ of finite index. For $i\in\{1,\dots,l\}$, write $d_{i}=(m_{i},m^{\prime}_{i})$ with $m_{i}\in M$ and $m_{i}^{\prime}\in M^{\prime}$. Note that the degrees $m_{i}$ (resp. $m^{\prime}_{i}$) generate a subgroup of $M$ (resp. $M^{\prime}$) of finite index. For $i\in\{1,\ldots,l\}$, write
@@
Let $M$ be a finitely generated abelian group, let $A,B$ be $M$-graded rings, and let $\Psi:A\to B$ be a morphism of $M$-graded rings.

### v5

@@ 1.1 Proj schemes
As part of his refoundation of Algebraic Geometry during the second part of the twentieth century, A. Grothendieck introduced the Proj scheme of an ${\mathbb{N}}$-graded ring. This generalization of the notion of projective space is natural in the framework of his theory of schemes. Grothendieck’s Proj construction has since been studied, taught, treated and used many thousands of times. During the first years of the third millennium, H. Brenner and S. Schröer generalized Grothendieck’s Proj construction to rings graded by arbitrary finitely generated abelian groups. Surprisingly, this very nice generalization still seems to be relatively confidential, although it has been recently used and studied in some references, including [Za19, KSU21, MRo24]. In this paper, we provide a slightly different perspective on the Brenner–Schröer Proj construction, and prove a couple of new results on this construction. In particular, we explain a generalization in this setting of Serre’s classical description of coherent sheaves on a projective space (or, more generally, a Proj scheme) as a Serre quotient of a category of graded modules over a graded ring, see e.g. [StP, Tag 01YR].
To illustrate this theory, we also explain how to describe flag varieties of reductive algebraic groups over algebraically closed fields, and vector bundles over such flag varieties associated with some representations of a Borel subgroup (including the Springer resolution) as the Proj scheme of a ring graded by characters of a maximal torus. The description of flag varieties as Proj schemes of $\mathbb{N}$-graded rings is classical, see e.g. [Wa], but it requires a choice of a strictly dominant weight. In the multi-graded setting one does not make any such choice, hence gets a more canonical construction, which is sometimes useful. In fact, versions of this construction appear implicitly or explicitly (but without proper references) in various constructions in Geometric Representation Theory, including [AB09, ABG04, ARd16], and it was one of our motivations for this work to make this construction completely explicit and rigorous. Note that [MS05, Exercise 14.16] indicates a special case of this construction.
Remark 1.1.
Other Proj-like constructions appear in the literature, often in relation with various forms of “toric geometry,” and some versions of the results explained below can be found in such frameworks (see e.g. [Roh14]). What we find particularly satisfying in the Brenner–Schröer theory is its elementary, while very general, nature.
@@ 1.2 The Brenner–Schröer construction
Let us now introduce the ideas of our work more precisely. Recall that given a commutative ring $A$, its (affine) spectrum ${\operatorname{Spec}}(A)$ is defined as the set of prime ideals of $A$, endowed with a certain topology and a sheaf of rings. Most treatments of Grothendieck’s Proj construction proceed by analogy with this construction, namely by associating to a graded ring the set of its graded prime ideals, and endowing it with a topology and a sheaf of rings after (somewhat tedious) identifications of various sets. However, Brenner–Schröer proceed differently in their generalization in [BS03, §2]: they start from the fact that, if $M$ is an abelian group, the datum of a grading on a commutative ring $A$ is equivalent to the datum of an action of the associated diagonalizable group scheme $\mathrm{D}_{{\operatorname{Spec}}(\mathbb{Z})}(M)$ on ${\operatorname{Spec}}(A)$ (see e.g. [SGA3, Exp. I, §4.7.3]), and then consider a certain open subset in a quotient ringed space constructed in terms of this action. Brenner–Schröer’s Proj scheme is covered by some special affine open subschemes, which generalize similar special affine subschemes appearing in Grothendieck’s Proj construction. In case $M=\mathbb{Z}$ and the negative-degree components in $A$ are 00, this recovers Grothendieck’s construction. (We will refer to this case as the “$\mathbb{N}$-graded case.”)
In the present treatment of the Proj construction, we put these “special affine subschemes” at the heart of the Proj construction. In fact, we give them a name: “potions.” Potions glue together to give birth to the Proj scheme thanks to a purely algebraic statement we call “the magic of potions.” To use this statement and glue potion schemes, we have to assume that $M$ is finitely generated. (In the $\mathbb{N}$-graded case, this result is part, as an auxiliary lemma, of the justification of the tedious identifications mentioned before in most treatments of the Proj construction.) Note that, in general, the underlying set of the Proj scheme is not the set of graded prime ideals of the given graded ring; it can be described as a set of graded ideals (see [BS03, Remark 2.3]), but we believe that this point of view is not really useful, and in any case it does not play any role in the present paper. Note that [MS05], already mentioned, provides some specific multi-graded Proj schemes (under the name spector) by glueing, cf. [MS05, Definition 10.25]. It seems that [BS03] and [MS05] are independent (each does not cite the other). Some other works related to multi-graded Proj schemes are cited in [BS03, Za19, KSU21, MRo24], see e.g. [Rob98, Pe07].
@@ 1.3 Contents
We now present the results contained in this document. Let as above $M$ be an abelian group and $A$ be a commutative ring endowed with a grading by $M$. In Sections 2 and 3 we define the notions of $M$-relevant families of $A$ (following Brenner–Schröer in the case of singletons, see Definition 2.3) and of potions associated with these families, and explain the definition of the Proj scheme ${\operatorname{Proj}}^{M}(A)$ of $A$ (in case $M$ is finitely generated). The potion associated with a relevant family $(f_{i}:i\in I)$ is the degree-00 part in the localization of $A$ with respect to the multiplicative subset generated by $(f_{i}:i\in I)$; see Definition 2.10. The spectra ${\operatorname{Spec}}(A_{(S)})$ are called potion schemes and are denoted $D_{\dagger}(S)$ (cf. Construction 3.1), and the scheme ${\operatorname{Proj}}^{M}(A)$ is defined by glueing the potions schemes $D_{\dagger}(S)$ over all relevant families (cf. Construction 3.1). This glueing is possible by the magic of potions (cf. Proposition 2.11). In other words, Section 2 focuses on commutative algebra around potions while Section 3 applies this material to explain the construction of Proj schemes.
In Sections 2 and 3, we also prove some basic results that may be known to experts. In particular, using the concept of quasi-relevant element (Definition 2.12), we prove that the Proj scheme of a tensor product of graded rings (with the product grading) is the fiber product of the Proj schemes of the rings (Proposition 3.16). We prove some compatibility results regarding the radical of the ideal generated by relevant elements (Propositions 2.6 and 2.8). In Example 2.14, we explain the relation between potions and dilatations of rings (as studied in [StP, MRR23, Ma24, Ma23]). We also show that the Proj construction is functorial (Proposition 3.13). In §3.6 we discuss multi-centered blowups and its relations with dilatations.
@@ 1.4 Relations between various constructions for graded rings
Recently, several other basic constructions for ${\mathbb{N}}$-graded rings were extended to more general gradings. In particular, in [Ma23a] the theory of ${\mathbb{G}}_{\mathrm{m}}$-attractors was refined and extended in the setting of actions of diagonalizable group schemes, and in [Ma24] mono-centered dilatations (involving ${\mathbb{N}}$-gradings) were extended to multi-centered dilatations. The algebraic point of view adopted in the present paper was partly inspired by these parallel generalizations. The diagram in Figure 1 summarizes some relations between these graded constructions. Lines materialize immediate connections. If two lines are parallel (not necessarily connected), then the associated connections are also parallel in some appropriate sense.
@@
  5. 5.
    Less recent and already widely used, the reference [HS04] provides another multi-graded generalization of a classical construction. It is possible that even more ${\mathbb{N}}$-graded constructions in schemes theory can be extended significantly to more general gradings.
Remark 1.2.
In relation with Comment (4) above, in Remark 3.4 we explain that one cannot refine the theory of schemes, in the framework of locally ringed spaces, so that ${\operatorname{Spec}}(S^{-1}A)\to{\operatorname{Spec}}(A)$ is an “open immersion” in this refined topology. Therefore, it seems really difficult geometrically to give meaning to ${\operatorname{Proj}}(A)$ if $A$ is graded by a non-finitely generated group in general. Note however that, algebraically, Proposition 2.11(2) holds for arbitrary graded rings.
@@ 1.5 Acknowledgements
We thank A.-M. Aubert, R. Bezrukavnikov, S. Brochard, J. P. dos Santos, A. Ducros, S. Evra, P. Scholze, T. Wedhorn and J. Zhang for useful discussions and/or comments related to the subject of this paper. We also thank the referee for very useful comments.
Foundational parts of this paper are formalized in Lean4 by A. Mayeux and J. Zhang [MZ25], in particular potions and Proj are implemented. Some of the conventions (e.g. the beginning of §1.6 and Definition 2.1) in the present paper were naturally adopted during the formalization in dependent type theory [MZ25].
@@ 1.6 Some notations and conventions
All the rings considered in this paper will be tacitly assumed to be associative and unital. (In practice, all the rings we want to consider will be commutative, but we will mention this assumption when it is necessary.)
Let us make explicit our conventions on graded rings. For this we fix an arbitrary abelian group $M$.
Definition 1.3.
An $M$-graded ring is a ring $A$ endowed with a direct sum decomposition $A=\bigoplus_{m\in M}A_{m}$ such that $A_{m}\cdot A_{m^{\prime}}\subset A_{m+m^{\prime}}$ for all $m,m^{\prime}\in M$.
Definition 1.4.
Let $A$ be an $M$-graded ring. A homogeneous pair in $A$ is a pair $(a,m)$ such that $a\in A$, $m\in M$ and $a\in A_{m}$. An element $a$ in $A$ is called homogeneous if there exists $m\in M$ such that $(a,m)$ is a homogeneous pair. The degree of a homogeneous pair $x=(a,m)$ is the element of $M$ defined as $\deg(x)=m$.
@@
Note that if $a$ is homogeneous and nonzero, there exists a unique $m$ such that $(a,m)$ is a homogeneous pair (because $A_{m}\cap A_{m^{\prime}}=\{0\}$ for $m\neq m^{\prime}$). For all $m\in M$, $(0,m)$ is a homogeneous pair.
Conventions ((Degrees of homogeneous elements)).
As a convention, given a homogeneous element $a$, the notation $\deg(a)$ means that we have fixed, implicitly or explicitly, an element $\deg(a)\in M$ such that $(a,\deg(a))$ is a homogeneous pair. The sentence “let $a$ be a homogeneous element of degree $m$” means “let $a$ be in $A_{m}$.” The element 00, regarded as an element of $A$, has no degree. But, as element of $A_{m}$ for some $m\in M$, 00 is an element of degree $m$.
Remark 1.6.
One can easily check that the unit $1\in A$ always belongs to $A_{0}$, in particular it is homogeneous.
If $R$ is a ring, an $M$-graded $R$-algebra is an $M$-graded ring $A$ endowed with a ring homomorphism $R\to A_{0}$. (In particular, $A$ is then an $R$-algebra.)
@@ 2.1 Graded rings and localizations
If $A$ is a commutative $M$-graded ring and $S\subset A$ is a homogeneous multiplicative subset, the localization $A_{S}$ of $A$ with respect to $S$ (which is denoted $S^{-1}A$ in many references, including [StP, Tag 00CM]) is canonically $M$-graded, explicitly, for $m\in M$ we have
$(A_{S})_{m}=\left\{\frac{a}{s}:\exists m^{\prime},m^{\prime\prime}\in M\text{ such that }a\in A_{m^{\prime}},\,s\in(S\cap A_{m^{\prime\prime}})\text{ and $m^{\prime}-m^{\prime\prime}=m$ }\right\}.$
Given a graded $A$-module $Q$, one can also consider the associated localization $Q_{S}$ (denoted $S^{-1}Q$ in [StP, Tag 07JZ]), which has a natural structure of a graded $A_{S}$-module.
@@ 2.2 Relevant families
Consider an abelian group $M$, and a commutative $M$-graded ring $A$. By a homogeneous subset of $A$ we mean a subset consisting of homogeneous elements.
@@
Let $S$ be a homogeneous subset of $A$. We denote by $\deg(S)$ the subset of $M$ defined as
$\deg(S):=\{m\in M:\exists s\in S,s\in A_{m}\}.$
Note that $\deg(\{0\})=M$. More generally, if $S$ is a homogeneous multiplicative subset of $A$, then $\deg(S)$ is a submonoid of $M$, also denoted $M[S\rangle$.
Definition 2.2.
Let $S$ be a homogeneous multiplicative subset of $A$.
@@
    Note that the homogeneous multiplicative subset $0^{\mathbb{N}}=\{0,1\}$ is always relevant.
@@
Proposition 2.5.
Let $S$ be a homogeneous multiplicative subset of $A$.
  1. 1.
    We have $M[\underline{S}]=\{m\in M:(A_{\underline{S}})^{\times}\cap(A_{\underline{S}})_{m}\neq\varnothing\}$, where $(A_{\underline{S}})^{\times}\subset A_{\underline{S}}$ is the group of invertible elements. In particular $S$ is $M$-relevant if and only if
    $M/\{m\in M:(A_{\underline{S}})^{\times}\cap(A_{\underline{S}})_{m}\neq\varnothing\}$
    is a torsion abelian group.
  2. 2.
    Assume that $M$ is finitely generated. A homogeneous element $x$ in $A$ is relevant if and only if there exist $k\in\mathbb{N}_{>0}$ and a factorization $x^{k}=x_{1}\cdots x_{l}$ into homogeneous factors $x_{i}$ of degree $d_{i}$ such that the degrees $d_{i}$ generate a subgroup of finite index in $M$.
Proof.
(1) The inclusion $M[\underline{S}]\subset\{m\in M:(A_{\underline{S}})^{\times}\cap(A_{\underline{S}})_{m}\neq\varnothing\}$ is obvious. Reciprocally let $m\in M$, and assume that there exists a fraction $\frac{a}{s}$ (with $a\in A$, $s\in\underline{S}$ homogeneous) which is invertible in $A_{\underline{S}}$ and of degree $m$. Since this element is invertible, there exist homogeneous elements $a^{\prime}\in A$ and $s^{\prime}\in\underline{S}$ such that $\frac{a}{s}\frac{a^{\prime}}{s^{\prime}}=1$. So there exists $s^{\prime\prime}\in\underline{S}$ such that $s^{\prime\prime}aa^{\prime}=ss^{\prime}s^{\prime\prime}$. This implies $a$ belongs to $\underline{S}$ and hence that $m=\deg(a)-\deg(s)$ belongs to $M[\underline{S}]$.
(2) Let $x$ be a relevant homogeneous element in $A$. Consider all homogeneous divisors of all positive powers of $x$, and the degrees of all these elements. By assumption, these degrees generate a subgroup $M^{\prime}$ of $M$ of finite index. Since $M$ is finitely generated, so is $M^{\prime}$ as a group. So we can find a finite number of homogeneous elements $x_{1},\dots,x_{m}$ such that their degrees generate $M^{\prime}$ and such that $x_{i}$ divides a power of $x$, say $x^{k_{i}}.$ Set $k=\sum_{k=1}^{m}k_{i}.$ Then the product $x_{1}\cdots x_{m}$ divides $x^{k}.$ So there exists $y$ such that $x^{k}=x_{1}\cdots x_{m}y.$ To finish the proof of the direct implication, it is enough to prove that $y$ can be chosen homogeneous, which follows from the following observation: if $x$ and $z$ are homogeneous in a graded ring and $z$ divides $x$, then there exists a homogeneous $y$ in $A$ such that $x=zy.$ The reverse implication is immediate. ∎
@@
Let $M$ and $M^{\prime}$ be two finitely generated abelian groups. Let $R$ be a commutative ring and let $A$ (resp. $A^{\prime}$) be a commutative $M$-graded (resp. $M^{\prime}$-graded) $R$-algebra. Then, considering $A\otimes_{R}A^{\prime}$ with its natural structure of a $M\times M^{\prime}$-graded $R$-algebra:
@@
(1) Let $x\in A\otimes_{R}A^{\prime}$ be relevant. By Proposition 2.5(2), there exist $k\in\mathbb{Z}_{>0}$ and a factorization $x^{k}=x_{1}\cdots x_{l}$ into homogeneous factors $x_{i}$ of degree $d_{i}\in M\times M^{\prime}$, such that the degrees $d_{i}$ generate a subgroup of $M\times M^{\prime}$ of finite index. For $i\in\{1,\dots,l\}$, write $d_{i}=(m_{i},m^{\prime}_{i})$ with $m_{i}\in M$ and $m_{i}^{\prime}\in M^{\prime}$. Note that the degrees $m_{i}$ (resp. $m^{\prime}_{i}$) generate a subgroup of $M$ (resp. $M^{\prime}$) of finite index. For $i\in\{1,\ldots,l\}$, write
@@
Let $M$ be a finitely generated abelian group, let $A,B$ be $M$-graded rings, and let $\Psi:A\to B$ be a homomorphism of $M$-graded rings.